# Zhicong Lin, Feihu Liu, Jiahang Liu§, Jing Liu, and Guoce Xin

### v1

@@ Abstract
Banderier, Marchal and Wallner considered Young tableaux with walls, which are similar to standard Young tableaux, except that local decreases are allowed at some walls. In this work, we prove a conjecture of Fuchs and Yu concerning the enumeration of two classes of three-row Young tableaux with walls. Combining with the work by Chang, Fuchs, Liu, Wallner and Yu leads to the verification of a conjecture on tree-child networks proposed by Pons and Batle. This conjecture was regarded as a specific and challenging problem in the Phylogenetics community until it was finally resolved by the present work.
@@ 1 Introduction
Young tableaux with walls were introduced by Banderier, Marchal and Wallner [1], and their enumeration is quite challenging. The main objective of this paper is to prove an enumerative conjecture (see Theorem 1.1) of Fuchs and Yu [9] concerning two classes of three-row Young tableaux with walls. Tree-child networks are one of the most prominent network classes for modeling evolutionary processes involving reticulation events. Our result together with the work by Chang, Fuchs, Liu, Wallner and Yu [8] leads to the verification of a conjecture due to Pons and Batle [18, Conjecture 1] on counting tree-child networks. Before proceeding to state our main result, we need to introduce some notations and definitions pertaining to Young tableaux with walls.
Let $\lambda=(\lambda_{1},\lambda_{2},\ldots)$ be a partition of $m$, denoted $\lambda\vdash m$. The Young diagram of $\lambda$ is obtained by drawing a collection of left-justified array of $m$ cells, where row $i$ has $\lambda_{i}$ cells reading form bottom to top. For the sake of convenience, the Young diagram of $\lambda$ is also denoted by $\lambda$. A standard Young tableau of $\lambda$ is an assignment of natural numbers $1,2,\ldots,m$ to the cells of the Young diagram, each number appearing once, such that rows are increasing from left to right, and columns are increasing from bottom to top.
Recently, Banderier, Marchal and Wallner [1, 2] considered a variation of standard Young tableaux, called Young tableaux with walls. For a Young diagram, we put a bold red edge (horizontally or vertically) between the consecutive cells which are allowed to be decreasing. Such places where a decrease is allowed (but not compulsory) are drawn by a bold red edge, which we call a “wall”. We refer to this type of Young diagram as Young diagram with walls; see Fig. 1 (a) for an example. As shown in Fig. 1 (b), a Young tableau with walls of $\lambda$ is a filling of the cells by labels $1,2,\dots,m=|\lambda|$, such that each label appears exactly once and the labels are increasing along each row and column (conditions for standard Young tableaux), but two labels separated by a wall need not be increasing.
@@
The enumeration of Young tableaux with walls is a difficult problem. Via bijections, hook-length-like formulas and the density method, some results for the enumeration of rectangular Young tableaux with walls were proved by Banderier, Marchal and Wallner [1]. Banderier and Wallner [2] further considered Young tableaux with periodic walls and reported more enumerative results and bijections with trees, lattice paths or permutations. By employing the kernel method and the lattice path enumeration theory, Liu and Xin [14] obtained the generating function for a class of two-row Young tableaux with periodic walls.
@@
A phylogenetic network is tree-child if, for each non-leaf node, at least one of its children is a tree node or leaf. See Fig. 4 for two examples of phylogenetic networks, where only the right one is tree-child. In a phylogenetic network with $n$ leaves, $k$ reticulation nodes and $t$ tree nodes, since the sum of the outdegrees equals that of indegrees, we have
@@
In fact, $a_{n,k}$ is interpreted in [18] as some Yamanouchi-like words, which are in one-to-one correspondence (see [23, Proposition 7.10.3 (d)]) with the Young tableaux with walls of shape $(n,n,k)$. Conjecture 1.2 stems form the work by Fuchs, Yu and Zhang [11] and has received considerable attention in recent years [8, 10]. In particular, Fuchs, Liu and Yu [10] confirmed Conjecture 1.2 for one-component tree-child networks which are an important building block of (general) tree-child networks (see [7]). Note that several interesting consequences of Conjecture 1.2 have been discussed in [18, 7].
@@ 2.2 An extension of $b_{n,k}$ and the introduction of its variant $\omega_{n,m,k}$
When $k=0$, the hook-length formula [23, Corollary 7.21.6] for counting standard Young tableau gives
@@ 2.3 Two key lemmas
This fact follows easily by linearity of the constant term operator. For rigorous definitions, see [24].
@@
Next, we eliminate $y$. Only the denominator factor $1+\frac{z}{4y(1+y)(1+z)}$ may results in negative powers of $y$. By quadratic formula, we can write
@@
Then apply Fact 2 with $u=Y_{1}$ gives
@@
Note that recurrence (17) for $\tilde{\omega}_{n,m,k}$ together with the initial conditions in (18) and (19) determines $\tilde{\omega}_{n,m,k}$ uniquely. To finish the proof, it remains to show that $\omega_{n,m,k}$ satisfies the same recurrence relation and initial conditions as that of $\tilde{\omega}_{n,m,k}$.
@@
2.5 Enumeration of tree-child networks
We now present several corollaries concerning the enumeration of tree-child networks. Pons and Batle [18, Proposition 8-12] derived the following results, Corollaries 2.13, 2.14, 2.15, 2.16, and 2.17 under the assumption that Conjecture 1.2 holds. We can now remove this assumption.
@@ 2.5 Enumeration of tree-child networks
where initial values $|\mathcal{TC}_{1,0}|=1$ and $|\mathcal{TC}_{i,-1}|=|\mathcal{TC}_{i,i}|=0$ for any $i$.
@@
By combining Equations (6) and (2), we obtain this result. ∎
@@
The cardinalities $|\mathcal{TC}_{n,k}|$ satisfy
@@
where initial values $|\mathcal{TC}_{1,0}|=1$ and $|\mathcal{TC}_{i,-1}|=|\mathcal{TC}_{i,i}|=0$ for any $i$.
@@
$a_{n,k+1}=\sum_{i=k+1}^{n}\prod_{j=i+1}^{n}(2j+k)a_{i,k}.$
The result follows from the relation above combined with Eq (6). ∎
@@
The cardinalities $|\mathcal{TC}_{n,k}|$ satisfy
@@
$a_{n+1,k}=\sum_{i=0}^{k}(2n+i+1)a_{n,i}.$
This result follows immediately from the above expression and Eq (6). ∎
@@
Let $\delta_{j}=j!\cdot\gamma_{j}$ for $j\geq 0$. The cardinalities $|\mathcal{TC}_{n,k}|$ satisfy
@@
where the coefficients $\delta_{i}$ are determined by the recursive equation
@@
This is a consequence of Theorem 2.1 together with Eq (6). ∎
@@
For $k$ fixed and $n\rightarrow\infty$, the cardinalities $|\mathcal{TC}_{n,k}|$ grow as
@@ 3.2 Functional equations for $B_{k}(x,t)$ via the kernel method
Using Theorem 3.2, we can easily compute expressions for $D_{k}(t)$ for $k\leq 140$ by Maple, which agree with the results from Theorem 3.1. This gives supports for the truth of Funchs and Yu’s enumerative conjecture (Theorem 1.1) in our earlier study.
@@ B.3 Characterization of $b_{n,k}$
By the definition of $e(D_{(n;i_{1},i_{2},\ldots,i_{k})})$ we see that this number is also equal to the number of labelings of the poset with integers $1,2,\ldots,2n+k$, such that every row and column is increasing along the direction of the arrow. Therefore, we obtain

### v2

@@ Abstract
Banderier, Marchal, and Wallner considered Young tableaux with walls, which are similar to standard Young tableaux, except that local decreases are allowed at some walls. In this work, we prove a conjecture of Fuchs and Yu concerning the enumeration of two classes of three-row Young tableaux with walls. Combining with the work by Chang, Fuchs, Liu, Wallner, and Yu leads to the verification of a conjecture on tree-child networks proposed by Pons and Batle. This conjecture was regarded as a specific and challenging problem in the Phylogenetics community until it was finally resolved by the present work.
@@ 1 Introduction
Young tableaux with walls were introduced by Banderier, Marchal, and Wallner [1], and their enumeration is quite challenging. The main objective of this paper is to prove an enumerative conjecture (see Theorem 1.1) of Fuchs and Yu [9] concerning two classes of three-row Young tableaux with walls. Tree-child networks are one of the most prominent network classes for modeling evolutionary processes involving reticulation events. Our result, together with the work by Chang, Fuchs, Liu, Wallner, and Yu [8], leads to the verification of a conjecture due to Pons and Batle [18, Conjecture 1] on counting tree-child networks. Before proceeding to state our main result, we need to introduce some notations and definitions related to Young tableaux with walls.
Let $\lambda=(\lambda_{1},\lambda_{2},\ldots)$ be a partition of $m$, denoted $\lambda\vdash m$. The Young diagram of $\lambda$ is obtained by drawing a collection of left-justified arrays of $m$ cells, where row $i$ has $\lambda_{i}$ cells reading from bottom to top. For the sake of convenience, the Young diagram of $\lambda$ is also denoted by $\lambda$. A standard Young tableau of $\lambda$ is an assignment of natural numbers $1,2,\ldots,m$ to the cells of the Young diagram, each number appearing once, such that rows are increasing from left to right, and columns are increasing from bottom to top.
Recently, Banderier, Marchal, and Wallner [1, 2] considered a variation of standard Young tableaux, called Young tableaux with walls. For a Young diagram, we put a bold red edge (horizontally or vertically) between the consecutive cells that are allowed to be decreasing. Such places where a decrease is allowed (but not compulsory) are drawn by a bold red edge, which we call a “wall”. We refer to this type of Young diagram as Young diagram with walls; see Fig. 1 (a) for an example. As shown in Fig. 1 (b), a Young tableau with walls of $\lambda$ is a filling of the cells by labels $1,2,\dots,m=|\lambda|$, such that each label appears exactly once and the labels are increasing along each row and column (conditions for standard Young tableaux), but two labels separated by a wall need not be increasing.
@@
The enumeration of Young tableaux with walls is a difficult problem. Via bijections, hook-length-like formulas, and the density method, some results for the enumeration of rectangular Young tableaux with walls were proved by Banderier, Marchal, and Wallner [1]. Banderier and Wallner [2] further considered Young tableaux with periodic walls and reported more enumerative results and bijections with trees, lattice paths, or permutations. By employing the kernel method and the lattice path enumeration theory, Liu and Xin [14] obtained the generating function for a class of two-row Young tableaux with periodic walls.
@@
A phylogenetic network is tree-child if, for each non-leaf node, at least one of its children is a tree node or leaf. See Fig. 4 for two examples of phylogenetic networks, where only the right one is tree-child. In a phylogenetic network with $n$ leaves, $k$ reticulation nodes, and $t$ tree nodes, since the sum of the outdegrees equals that of the indegrees, we have
@@
In fact, $a_{n,k}$ is interpreted in [18] as some Yamanouchi-like words, which are in one-to-one correspondence (see [23, Proposition 7.10.3 (d)]) with the Young tableaux with walls of shape $(n,n,k)$. Conjecture 1.2 stems from the work by Fuchs, Yu, and Zhang [11] and has received considerable attention in recent years [8, 10]. In particular, Fuchs, Liu, and Yu [10] confirmed Conjecture 1.2 for one-component tree-child networks which are an important building block of (general) tree-child networks (see [7]). Note that several interesting consequences of Conjecture 1.2 have been discussed in [18, 7]; see also Sec. 2.5 for some of them.
@@ 2.2 An extension of $b_{n,k}$ and the introduction of its variant $\omega_{n,m,k}$
When $k=0$, the hook-length formula [23, Corollary 7.21.6] for counting standard Young tableaux gives
@@ 2.3 Two key lemmas
This fact follows easily from the linearity of the constant term operator. For rigorous definitions, see [24].
@@
Next, we eliminate $y$. Only the denominator factor $1+\frac{z}{4y(1+y)(1+z)}$ may results in negative powers of $y$. By the quadratic formula, we can write
@@
Then applying Fact 2 with $u=Y_{1}$ gives
@@
Note that recurrence (17) for $\tilde{\omega}_{n,m,k}$ together with the initial conditions in (18) and (19) determines $\tilde{\omega}_{n,m,k}$ uniquely. To finish the proof, it remains to show that $\omega_{n,m,k}$ satisfies the same recurrence relation and initial conditions as those of $\tilde{\omega}_{n,m,k}$.
@@
2.5 Applications to the enumeration of tree-child networks
We now present several corollaries concerning the enumeration of tree-child networks. Pons and Batle [18, Propositions 8-12] derived the following results, Corollaries 2.13, 2.14, 2.15, 2.16, and 2.17 under the assumption that Conjecture 1.2 holds. We can now remove this assumption.
@@ 2.5 Applications to the enumeration of tree-child networks
where the initial values are $|\mathcal{TC}_{1,0}|=1$ and $|\mathcal{TC}_{i,-1}|=|\mathcal{TC}_{i,i}|=0$ for any $i$.
@@
By combining Eqs. (6) and (2), we obtain this result. ∎
@@
The sequence $|\mathcal{TC}_{n,k}|$ satisfies
@@
where the initial values are $|\mathcal{TC}_{1,0}|=1$ and $|\mathcal{TC}_{i,-1}|=|\mathcal{TC}_{i,i}|=0$ for any $i$.
@@
$a_{n+1,k}=\sum_{i=0}^{k}(2n+i+1)a_{n,i}.$
The result then follows immediately from the relation above combined with Eq. (6). ∎
@@
The sequence $|\mathcal{TC}_{n,k}|$ satisfies
@@
$a_{n,k+1}=\sum_{i=k+1}^{n}\prod_{j=i+1}^{n}(2j+k)a_{i,k}.$
This result follows from the above expression and Eq. (6). ∎
@@
Let $\delta_{j}=j!\cdot\gamma_{j}$ for $j\geq 0$. The sequence $|\mathcal{TC}_{n,k}|$ satisfies
@@
where the coefficients $\delta_{i}$ are determined by the recursion
@@
This is a consequence of Theorem 2.1 together with Eq. (6). ∎
@@
For $k$ fixed and $n\rightarrow\infty$, the sequence $|\mathcal{TC}_{n,k}|$ grows as
@@ 3.2 Functional equations for $B_{k}(x,t)$ via the kernel method
Using Theorem 3.2, we can easily compute expressions for $D_{k}(t)$ for $k\leq 140$ by Maple, which agree with the results from Theorem 3.1. This gives support for the truth of Funchs and Yu’s enumerative conjecture (Theorem 1.1) in our earlier study.
@@ B.3 Characterization of $b_{n,k}$
By the definition of $e(D_{(n;i_{1},i_{2},\ldots,i_{k})})$, we see that this number is also equal to the number of labelings of the poset with integers $1,2,\ldots,2n+k$, such that every row and column is increasing along the direction of the arrow. Therefore, we obtain